OpenStudy (anonymous):

Solve 2 log x = log 36 Solve 2 log x = log 36 @Mathematics

OpenStudy (anonymous):

\[\log(x^2)=\log(36)\] \[x^2=36\] \[x=6\]

OpenStudy (anonymous):

It would be + / - 6?

OpenStudy (anonymous):

*wouldnt

OpenStudy (anonymous):

GOOD QUESTION! actually to solve \[x^2=36\] you are supposed to say \[x=\pm6\] but you can ignore the negative answer because you cannot take the log of a negative number

OpenStudy (anonymous):

Oh ok, so what would this property be known as?